# Discovering Artificial Viscosity Models for Discontinuous Galerkin approximation of Conservation Laws using Physics-Informed Machine Learning

### v1

@@ 3.4 The neural viscosity approach
The way neural viscosity works is macroscopically the same in which any AV model works (as described in Section 3.1). The NN works independently on each cell, it receives as inputs some local quantities, and it outputs an artificial viscosity value $\mu_{K}^{\text{bnd}}$ for each cell $K$ and finally, it performs a global smoothing step. Let us now describe this procedure in detail. The first difference with respect to the original neural viscosity model is that, instead of building and training a different NN for each polynomial degree $k>0$, we build just one independently of $k$, reducing training costs. This is enabled by the fact that instead of feeding the local nodal values of the solution $\boldsymbol{U}^{n}|_{K}$ (the evaluation of $u_{h}$ at time $t^{n}$ in the degrees of freedom of $K$, which is equal to $u_{h}|_{K}(t^{n})$ since we are employing a nodal basis) to the NN, we extract some representative statistics of these values, namely the mean and the standard deviation and we add as input also the local polynomial degree $k$. The input can be enriched with other information such as the median, the quartiles, the minimum, or the maximum. In Section 4.3 we present numerical experiments that show the tradeoff between accuracy and computational cost of this choice. The second difference is that to accelerate the training, we feed not only the current velocity but also the nodal values of
  • the gradient of the velocity;
  • the jump of the velocity across the element boundary, since it is a key information to assess the smoothness of the current solution;
  • the velocity at the previous time step, to provide information about temporal correlation;
  • flux of the velocity, to provide information about the physics of the problem.
@@
Observe that we denote with $y$ the output of the NN since it is a scalar quantity. Moreover, the scaling is time dependent, so it changes at every time step. The insight behind this rescaling is that $\Lambda$ makes the same network work for difference PDEs since it represents the maximum local wave speed, and $\tilde{h}$ provides the optimal viscosity scaling $h^{k+1}$ in the presence of smooth solution [31]. Indeed, even if the NN prescribes low dissipation when the solution is regular, due to the nature of the softplus function, it is never zero, which prevents the method from having optimal convergence. Instead, $\tilde{h}$ is a parameter-free option that provides the following scaling
@@
which enables a correct reduction of the artificial viscosity when the solution is smooth. A complete scheme of what we have just described is shown in Figure 1.
@@ 4 Physics-informed machine learning
In this section, we describe in detail how the neural network is trained. We propose a novel approach that is inspired by reinforcement learning and is enhanced by physics-informed machine learning. Namely, thanks to automatic differentiation, we can differentiate through the DG solver (the environment in RL terminology) and define a loss function (the opposite of the reward) as the norm of the error between a reference solution and the solution obtained with the current parameters of the NN (the policy) that surrogates a viscosity model (agent). The loss is also enhanced with suitable regularization terms. To conclude the comparison with RL terminology, the local statistics extracted from each mesh element are the state and the viscosity $\mu_{K}^{\text{bnd}}$ the action the agent chooses. The major difference with respect to RL is that we are able to differentiate through the environment. Thus, by using a gradient-based optimizer, we can directly compute updates to the parameter of the NN so to minimize the loss (maximize the reward). Hence, this training algorithm enables us to discover a new neural viscosity model that has the property of minimizing the norm of the error. This approach allows us to avoid building a dataset with a reference value of artificial viscosity. Indeed, many RL algorithms overcome the problem of a not differentiable environment using a second neural network called critic that surrogates it. This is however expensive, since the environment dynamic must also be learnt, and may lead to instabilities due to the interaction of the policy with the critic [39].
@@ 4.1 Loss function
where the variable $\boldsymbol{\theta}$ explicits the dependence on the parameters of the NN. Hence, given a discrete reference solution $\boldsymbol{U}_{ex}^{n}$, that is the evaluation of a reference solution $\boldsymbol{u}_{ex}:\Omega\times(0,T]\to\mathbb{R}^{m}$ at the degrees of freedom at time $t^{n}$, by employ automatic differentiation we can compute
@@
$\frac{\partial}{\partial\boldsymbol{\theta}}\left\|\boldsymbol{U}^{n}(\boldsymbol{U}^{0};\boldsymbol{\theta})-\boldsymbol{U}_{ex}^{n}\right\|_{p}\quad\forall\>n>0.$
@@
This is one of the main ingredients of our loss: using an optimizer, we can find the set of parameter $\boldsymbol{\theta}$ that defines the neural viscosity model that minimizes the error with respect to the reference solution. We define the $p$-loss as follows:
@@
$\mathcal{L}(\boldsymbol{\theta};\mathscr{P})=\sum_{n=1}^{N}\left[\left\|\boldsymbol{U}^{n}(\boldsymbol{\theta})-\boldsymbol{U}_{ex}^{n}\right\|_{p}^{p}+w_{0}\left\|\nabla\boldsymbol{U}^{n}(\boldsymbol{\theta})-\nabla\boldsymbol{U}_{ex}^{n}\right\|_{p}^{p}+\sum_{\ell_{r,p}\in\mathcal{R}}w_{r}\ell_{r,p}^{p}(\boldsymbol{U}^{n}(\boldsymbol{\theta}),\boldsymbol{U}_{ex}^{n})\right],$ (14)
where $\mathcal{R}$ is a set of regularization terms $\ell_{r,p}$ in $p$-norm and
$\mathscr{P}=\{\mathcal{K}_{h},k,N,\boldsymbol{f},\Delta t,\boldsymbol{U}^{0},\boldsymbol{u}_{ex}\},$ (15)
@@
is a tuple that uniquely defines a problem. Namely, a triangulation $\mathcal{K}_{h}$, polynomial degree of the basis $k$, number of timesteps $N$, physical flux $\boldsymbol{f}$, time discretization $\Delta t$, initial condition $\boldsymbol{U}^{0}$ and the exact solution $\boldsymbol{u}_{ex}$. Notice that, for the sake of simplicity, we have removed the explicit dependence on $\boldsymbol{U}^{0}$ of $\boldsymbol{U}^{n}$ for $n>0$. Indeed, in this loss function, it is implicit the fact that the DG solver described in Section 2.2 is employed to compute $\boldsymbol{U}^{n}$ from $\boldsymbol{U}^{0}$. Notice that with a slight abuse of notation, we apply (linear) differential and integral operators to $\boldsymbol{U}^{n}$, with the meaning of considering the vector of coefficients describing the functions obtained by applying said operator to $\boldsymbol{u}_{h}(\boldsymbol{x},t)$. The metrics that we consider are the following:
@@
    $\epsilon(t^{n})=\left\|\boldsymbol{U}^{n}-\boldsymbol{U}_{ex}^{n}\right\|_{p}.$ (16)
@@
    $\nabla{\epsilon}(t^{n})=\left\|\nabla\boldsymbol{U}^{n}-\nabla\boldsymbol{U}_{ex}^{n}\right\|_{p}.$ (17)
@@
    $[\![\epsilon]\!](t^{n})=\left\|[\![\boldsymbol{U}^{n}-\boldsymbol{U}_{ex}^{n}]\!]\right\|_{p}.$ (18)
@@
    where $\mathbbm{1}_{S}$ is the indicator function of a set $S$. The amplitude of overshoots and undershoots with respect to a reference solution is one of the main criteria with which to evaluate a numerical solution since it provides insights into the behavior and stability of a system.
@@
    $\text{vp}(t^{n})=\frac{\frac{1}{\left|K\right|}\int_{K}\mu}{\left\|\nabla\boldsymbol{U}^{n}\right\|_{p}+\varepsilon},$ (21)
@@
Not all of them are used in the final version of the loss, indeed we use the minimal number of terms that produce a good model since each term has a weight $w_{r}$ that needs to be tuned (just once offline). The other terms are still used as metrics to assure the quality of the neural viscosity model. Namely, we have found that the best approach is to employ the overshoot/undershoot and the viscosity penalization.
Another key choice is the value of $p$. We pick $p=1$ since it tends to emphasize sparsity in the data, meaning it is less sensitive to small values and can be more robust in the presence of outliers. Namely, it is more appropriate in the presence of oscillations that have sparse and isolated peaks, which is our case.
@@ 4.2 Training algorithm
1:procedure Training($\{\mathscr{P}_{j}\}_{j=1}^{N_{p}}$, $\{\mathscr{P}^{\text{test}}\}_{j=1}^{N_{pt}}$, lr, scheduler, $N_{E}$, $N_{B}$, $\boldsymbol{\theta}$, $\mathcal{L}$, $L_{\max{}}$, $n_{\textnormal{test}}$)
@@
10:                  $L\leftarrow L+\mathcal{L}(\boldsymbol{\theta};\mathscr{P}_{j})$triangleright $\mathcal{L}$ is defined by $p$ and the weights $\{w_{i}\}_{i\in\mathcal{R}}$, see Eq. (14)
@@
Let us now detail the novel algorithm through which the training is performed. In particular, this training process does not belong to the paradigm of supervised learning since the explicit value of viscosity that the neural network should learn for a certain value of inputs is never computed. The training algorithm is stated in Algorithm 1 with detailed comments and is summarized in Figure 2. It works as follows. It takes as inputs:
@@
  • A learning rate scheduler. We employ a simple “reduce on plateaus” [47] with patience 30 and reduction factor equal to $\frac{1}{2}$.
@@
  • $\mathcal{L}$, the loss function defined in Eq. (14). It is defined by $p$ and the weights $\{w_{i}\}_{i\in\mathcal{R}}$.
@@
Let us now briefly discuss the choice of hyperparameters for this training algorithm. We observe that we obtain better results with a small batch size, usually between two and four. A batch size equal to one is also a viable option, however, we notice that if the batch size is too large ($N_{B}>32$) the algorithm stops making progress in the loss minimization. This issue is either related to the optimizer or numerical instability in the AD. Moreover, the batches are shuffled between each epoch in order to increase the stochasticity.
In this training algorithm, it is paramount to choose $N$ properly: if it is too small, the neural network does not have enough data to generalize, if it is too large we may incur into exploding gradient problems unless the neural network has a proper weight initialization. We propose a multi-step pre-train in which we iterate the training Algorithm 1 for increasing values of $N$ for a few epochs, starting from $N=N_{0}$ up to the desired final $N$ with a certain step $N_{e}$. The increase in number of timesteps $N_{e}$ is computed adaptively, namely $N_{e}=\texttt{clip}(N_{e}^{\min{}},N_{e}^{\alpha},N_{e}^{\max{}})$, where $N_{e}^{\alpha}$ is the largest integer so that the average loss per timestep with $N+N_{e}^{\alpha}$ timesteps is less than a constant (hyperparameter) $\alpha_{e}>0$ multiplied by the average loss per timestep with $N$ timesteps. Once the NN reaches good generalization properties it is observed that the average loss per timestep decreases when $N$ increases. Indeed, this shows that the neural viscosity model is making accurate predictions of the “future” (data it has never seen before). Figure 3 shows an example of the training loss. This process is expensive from a computational point of view, however, it enables the NN to discover a viscosity model without any biases.

### v2

@@ 3.4 The neural viscosity approach
The way neural viscosity works is macroscopically the same in which any AV model works (as described in Section 3.1). The NN works independently on each cell, it receives as inputs some local quantities, and it outputs an artificial viscosity value $\mu_{K}^{\text{bnd}}$ for each cell $K$ and finally, it performs a global smoothing step. Let us now describe this procedure in detail. The first difference with respect to the original neural viscosity model is that, instead of building and training a different NN for each polynomial degree $k>0$, we build just one independently of $k$, reducing training costs. This is enabled by the fact that instead of feeding the local nodal values of the solution $\boldsymbol{U}^{n}|_{K}$ (the evaluation of $u_{h}$ at time $t^{n}$ in the degrees of freedom of $K$, which is equal to $u_{h}|_{K}(t^{n})$ since we are employing a nodal basis) to the NN, we extract some representative statistics of these values, namely the mean and the standard deviation and we add as input also the local polynomial degree $k$. The input can be enriched with other information such as the median, the quartiles, the minimum, or the maximum. In Section 4.3 we present numerical experiments that show the tradeoff between accuracy and computational cost of this choice. The second difference is that to accelerate the training, we employ as inputs of the neural network not only the current state variable but all the following features:
  • state variable;
  • the gradient of the state variable;
  • the jump of the state variable across the element boundary, since it is a key information to assess the smoothness of the current solution;
  • the state variable at the previous time step, to provide information about temporal correlation;
  • flux of the state variable, to provide information about the physics of the problem.
@@
Observe that we denote with $y$ the output of the NN since it is a scalar quantity. Moreover, the scaling is time dependent, so it changes at every time step. The insight behind this rescaling is that $\Lambda$ makes the same network work for different PDEs since it represents the maximum local wave speed, and $\tilde{h}$ provides the optimal viscosity scaling $h^{k+1}$ in the presence of smooth solution [31]. Indeed, even if the NN prescribes low dissipation when the solution is regular, due to the nature of the softplus function, it is never zero, which prevents the method from having optimal convergence. Instead, $\tilde{h}$ is a parameter-free option that provides the following scaling
@@
which enables a correct reduction of the artificial viscosity when the solution is smooth. In other words, at each timestep, the neural network prescribes a single viscosity value for each element of the mesh. A complete scheme of what we have just described is shown in Figure 1.
@@ 4 Physics-informed machine learning
In this section, we describe in detail how the neural network is trained. We propose a novel approach that is inspired by reinforcement learning and is enhanced by physics-informed machine learning. Namely, thanks to automatic differentiation, we can differentiate through the DG solver (the environment in RL terminology) and define a loss function (the opposite of the reward) as the norm of the error between a reference solution and the solution obtained with the current parameters of the NN (the policy) that surrogates a viscosity model (agent). The loss is also enhanced with suitable regularization terms. To conclude the comparison with RL terminology, the local statistics extracted from each mesh element are the state and the viscosity $\mu_{K}^{\text{bnd}}$ the action the agent chooses. The major difference with respect to RL is that we are able to differentiate through the environment. Thus, by using a gradient-based optimizer, we can directly compute updates to the parameter of the NN so to minimize the loss (maximize the reward). Hence, this training algorithm enables us to discover a new neural viscosity model that has the property of minimizing the norm of the error. This approach allows us to avoid building a dataset with a reference value of artificial viscosity. Indeed, many RL algorithms overcome the problem of a non-differentiable environment using a second neural network called critic that surrogates it. This is however expensive, since the environment dynamic must also be learnt, and may lead to instabilities due to the interaction of the policy with the critic [39]. The choice of using physics-informed machine learning, and in particular automatic differentiation through the DG solver, is motivated mainly by efficiency. Indeed, automatic differentiation enables the integration of the DG solver dynamics into neural network training, ensuring precision and efficiency. Concerning the architecture, the specific use of a dense feed-forward neural network is motivated by its simplicity, flexibility, and universality. They are straightforward to implement, computationally efficient, and can be tailored to various problems by adjusting their architecture.
@@ 4.1 Loss function
where the variable $\boldsymbol{\theta}$ explicits the dependence on the parameters of the NN. Hence, given a discrete reference solution $\boldsymbol{U}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ref}}^{n}$, that is the evaluation of a reference solution $\boldsymbol{u}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ref}}:\Omega\times(0,T]\to\mathbb{R}^{m}$ at the degrees of freedom at time $t^{n}$, by employ automatic differentiation we can compute
@@
$\frac{\partial}{\partial\boldsymbol{\theta}}\left\|\boldsymbol{U}^{n}(\boldsymbol{U}^{0};\boldsymbol{\theta})-\boldsymbol{U}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ref}}^{n}\right\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}}\quad\forall\>n>0.$
@@
This is one of the main ingredients of our loss: using an optimizer, we can find the set of parameter $\boldsymbol{\theta}$ that defines the neural viscosity model that minimizes the error with respect to the reference solution. More precisely, the solution $\boldsymbol{u}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ref}}$ is either the analytical solution (when available) or an overkill solution obtained using the DG solver on a triangulation that has a mesh size $h$ that is sufficiently small (in our computations eight times smaller). We define the ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}$-loss as follows:
@@
$\mathcal{L}(\boldsymbol{\theta};\mathscr{P})=\sum_{n=1}^{N}$ $\bigg{[}\left\|\boldsymbol{U}^{n}(\boldsymbol{\theta})-\boldsymbol{U}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ref}}^{n}\right\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}}+w_{0}\left\|\nabla\boldsymbol{U}^{n}(\boldsymbol{\theta})-\nabla\boldsymbol{U}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ref}}^{n}\right\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}}$ (15)
$+\sum_{\ell_{r,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}}\in\mathcal{R}}w_{r}\ell_{r,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}}(\boldsymbol{U}^{n}(\boldsymbol{\theta}),\boldsymbol{U}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ref}}^{n})\bigg{]},$
where $\mathcal{R}$ is a set of regularization terms $\ell_{r,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}}$ in ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}$-norm that will be defined below and
$\mathscr{P}=\{\mathcal{K}_{h},k,N,\boldsymbol{f},\Delta t,\boldsymbol{U}^{0},\boldsymbol{u}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ref}}\},$ (16)
@@
is a tuple that uniquely defines a problem. Namely, a triangulation $\mathcal{K}_{h}$, polynomial degree of the basis $k$, number of timesteps $N$, physical flux $\boldsymbol{f}$, time discretization $\Delta t$, initial condition $\boldsymbol{U}^{0}$ and the reference solution $\boldsymbol{u}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ref}}$. Notice that, for the sake of simplicity, we have removed the explicit dependence on $\boldsymbol{U}^{0}$ of $\boldsymbol{U}^{n}$ for $n>0$. Indeed, in this loss function, it is implicit the fact that the DG solver described in Section 2.2 is employed to compute $\boldsymbol{U}^{n}$ from $\boldsymbol{U}^{0}$. Notice that with a slight abuse of notation, we apply (linear) differential and integral operators to $\boldsymbol{U}^{n}$, with the meaning of considering the vector of coefficients describing the functions obtained by applying said operator to $\boldsymbol{u}_{h}(\boldsymbol{x},t)$. The metrics that we consider (which is a superset of the regularization terms $\ell_{r,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}}$ in $\mathcal{R}$) are the following:
@@
    $\epsilon(t^{n})=\left\|\boldsymbol{U}^{n}-\boldsymbol{U}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ref}}^{n}\right\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}}.$ (17)
@@
    $\nabla{\epsilon}(t^{n})=\left\|\nabla\boldsymbol{U}^{n}-\nabla\boldsymbol{U}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ref}}^{n}\right\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}}.$ (18)
@@
    $[\![\epsilon]\!](t^{n})=\left\|[\![\boldsymbol{U}^{n}-\boldsymbol{U}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ref}}^{n}]\!]\right\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}}.$ (19)
@@
    $+\left\|\boldsymbol{U}^{n}\mathbbm{1}_{\boldsymbol{U}^{n}<\min\boldsymbol{U}^{n}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ref}}}-\min\boldsymbol{U}^{n}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ref}}\right\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}},$
    where $\mathbbm{1}_{S}$ is the indicator function of a set $S$. The amplitude of overshoots and undershoots with respect to a reference solution is one of the main criteria with which to evaluate a numerical solution since it provides insights into the behavior and stability of a system. To apply backpropagation to this term, we do not implement it as a product of an indicator function, but we restrict the computations only to the entries of $\boldsymbol{U}^{n}$ that meet the condition stated in the indicator function.
@@
    $\text{vp}(t^{n})=\frac{\frac{1}{\left|K\right|}\int_{K}\mu}{\left\|\nabla\boldsymbol{U}^{n}\right\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}}+\varepsilon},$ (22)
@@
Not all of them are used in the final version of the loss, indeed we use the minimal number of terms that produce a good model since each term has a weight $w_{r}$ that needs to be tuned (just once offline). The other terms are still used as metrics to assure the quality of the neural viscosity model. Namely, we have found that the best approach is to employ as the set of regularization terms $\ell_{r,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}}$ in $\mathcal{R}$ the overshoot/undershoot (o/u) and the viscosity penalization (vp). The weights in the loss function for these two terms are hyperparameters to tune.
Another key choice is the value of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}$. We pick ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}=1$ since it tends to emphasize sparsity in the data, meaning it is less sensitive to small values and can be more robust in the presence of outliers. Namely, it is more appropriate in the presence of oscillations that have sparse and isolated peaks, which is our case. Indeed, we have tested other values of $q$, such as $q=2$, and they produced more oscillating results.
@@ 4.2 Training algorithm
1:procedure Training($\{\mathscr{P}_{j}\}_{j=1}^{N_{p}}$, $\{\mathscr{P}_{j}^{\text{test}}\}_{j=1}^{N_{pt}}$, lr, scheduler, $N_{E}$, $N_{B}$, $\boldsymbol{\theta}$, $\mathcal{L}$, $L_{\max{}}$, $n_{\textnormal{test}}$)
@@
10:                  $L\leftarrow L+\mathcal{L}(\boldsymbol{\theta};\mathscr{P}_{j})$triangleright $\mathcal{L}$ is defined by $q$ and the weights $\{w_{i}\}_{i\in\mathcal{R}}$, see Eq. (15)
@@
Let us now detail the novel algorithm through which the training is performed. In particular, this training process does not belong to the paradigm of supervised learning since the explicit value of viscosity that the neural network should learn for a certain value of inputs is never computed. Our approach is built on the idea that target viscosity values are not directly used to train the model. Instead, we utilize reference solutions to guide the learning process. However, the approach can neither be classified as unsupervised: when an exact solution is not available, a reference solution must be built using a viscosity model with suitably chosen parameters. The training algorithm is stated in Algorithm 1 with detailed comments and is summarized in Figure 2. It works as follows. It takes as inputs:
@@
  • A learning rate scheduler [40]. We employ a simple “reduce on plateaus” [47] with patience 30 (number of epochs to wait for a loss reduction) and reduction factor equal to $\frac{1}{2}$ (factor by which the learning rate is reduced).
@@
  • $\mathcal{L}$, the loss function defined in Eq. (15). It is defined by ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q}$ and the weights $\{w_{i}\}_{i\in\mathcal{R}}$.
@@
Let us now briefly discuss the choice of hyperparameters for this training algorithm. We observe that we obtain better results with a small batch size, usually between two and four. It is paramount to choose $N$ properly: if it is too small, the neural network does not have enough data to generalize, if it is too large we may incur into exploding gradient problems unless the neural network has a proper weight initialization. We propose a multi-step pre-train in which we iterate the training Algorithm 1 for increasing values of $N$ for a few epochs, starting from $N=N_{0}$ up to the desired final $N$ with a certain step $N_{e}$. The increase in number of timesteps $N_{e}$ is computed adaptively, namely, defined
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\texttt{clip}(a,b,c)=\max(\min(b,c),a),}$
we compute $N_{e}$ as $N_{e}=\texttt{clip}(N_{e}^{\min{}},N_{e}^{\alpha},N_{e}^{\max{}})$, where $N_{e}^{\alpha}$ is the largest integer so that the average loss per timestep with $N+N_{e}^{\alpha}$ timesteps is less than a constant (hyperparameter) $\alpha_{e}>0$ multiplied by the average loss per timestep with $N$ timesteps. Once the NN reaches good generalization properties it is observed that the average loss per timestep decreases when $N$ increases. Indeed, this shows that the neural viscosity model is making accurate predictions of the “future” (data it has never seen before). Figure 3 shows an example of the training loss. This process is expensive from a computational point of view, however, it enables the NN to discover a viscosity model without any biases.